OpenStudy (anonymous):

evaluate abc over 2d if a=-4,b=3,c=-7 and d=6

TheSmartOne (thesmartone):

\[\frac{ abc }{ 2d }\]

TheSmartOne (thesmartone):

\[\frac{ (-4)(3)(-7) }{ 2(6) }\]

TheSmartOne (thesmartone):

Can you do it from ther @anam123

TheSmartOne (thesmartone):

there*

OpenStudy (confluxepic):

Just plug in the values and solve.

OpenStudy (anonymous):

i will give a medal

TheSmartOne (thesmartone):

-4*-7=+28 and 6*2=12 \[\frac{ 28\times3 }{ 12 }\]

OpenStudy (anonymous):

14

TheSmartOne (thesmartone):

28*3=84 84/12=7 So \(\Huge\color{blue}{\frac{abc}{2d}=7}\)

OpenStudy (anonymous):